the cost in dollars of manufacturing w wingnuts is given by the expression 0.07w and 13.3. find the cost when 375 wing nuts are made. what is the average cost to manufacture one wing nut?
Answer by tutorcecilia(2152)

c = 0.07w + 13.3
c = 0.07w + 13.3 [Plug-in x=375 and solve for "c"]
c = 0.07(375)
c=39.55
.
Check by plugging (c=39.55) back into the original equation and solve for "w"
c = 0.07w + 13.3
39.55 = 0.07w + 13.3
39.55 -13.33 = 0.07w + 13.3 - 13.33
26.22/.07=.07w/.07
374.57=w [Round off to 375]
.
For one wignut, plug-in x=1 and solve for "c":
c = 0.07w + 13.3
c = 0.07(1) + 13.3
c = 0.07(1) + 13.3
c=13.37
.
Checking:
13.37=0.07(1) + 13.3
13.37=13.37 [Checks out]

w + 13.3
. find the cost when 375 wing nuts are made. what is the average cost to manufacture one wing nut?